OpenStudy (anonymous):

Solve the system of linear equations below. 2x + 3y = 2 x + 6y = 4

OpenStudy (anonymous):

/| 2*x+3*y-2 = 0| x+6*y-4 = 0 We try to solve the equation: 2*x+3*y-2 = 0 2*x+3*y-2 = 0 // - 2*x-2 3*y = -(2*x-2) // : 3 y = (-(2*x-2))/3 y = (2-(2*x))/3 We insert the solution into one of the initial equations of our system of equations We get a system of equations: /| 6*((2-(2*x))/3)+x-4 = 0| y = (2-(2*x))/3 (6*(2-2*x))/3+x-4 = 0 4-3*x-4 = 0 -3*x = 0 -3*x = 0 // : -3 x = 0 We insert the solution into one of the initial equations of our system of equations For y = (2-(2*x))/3: y = (2-(0*2))/3 y = (0+2)/3 y = 2/3 We get a system of equations: /| y = 2/3| x = 0

OpenStudy (anonymous):

do you have any question about that?

OpenStudy (anonymous):

you can also use elimination method

OpenStudy (anonymous):

if you have any question you can massage me back.

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

hola como esestas

OpenStudy (anonymous):

x = 0, y = 2 3 x = 2, y = - 2 3 x = 4 3 , y = 2 9 x = 4, y = -2

OpenStudy (anonymous):

@joojoe those the answer you can choose from